# Survey of Graph Neural Network for Internet of Things and NextG Networks

### v1

@@ VII Opportunities and Roadmap ahead
As seen in the article, GNNs have been widely adopted for a wide range of applications but several areas are in search of breakthroughs to make GNNs more impactful. This provides an opportunity for the research community to further develop new features for GNN as well as apply GNN for even broader research areas. To this end, in this section, we discuss the possible avenues of research to improve the capabilities of GNN and research areas where GNN will play a crucial role in IoT and NextG networks. Figure 14 shows the key areas of research opportunities.
@@
    Scalable and Reliable GNN. Scalability is one of the crucial aspects to gauge the success of any ML algorithm and GNN is not an exception. As of now, most of the GNN algorithms are tested on a smaller network topology and present good results with a slightly larger network to show the scalability. However, with the evolving number of IoT devices and wireless devices, it is crucial to develop algorithms that can scale up to thousands or hundreds of thousands of devices without any possible degradation in performance. GNN promises to be a great tool that can enable scalable ML models, it is important for the research community to further investigate and rigorously test their algorithm from the perspective of scalability on real networks.
    Reliability is another important area that requires further attention. As of now, most of the GNN-based algorithms designed assume a perfect environment with pre-defined interference, channel, and other relevant models for system design. It is important to design GNNs that are robust to adversarial attacks, sensor malfunction, data corruption, and other unforeseen conditions. This is especially important when there are millions of connected devices and data from one or more sensors should not impact the whole network. This can be addressed by systematically identifying, quantifying, and rectifying anomalies or security risks. This promises to be an interesting area of research and will significantly benefit in maturation and adoption of GNN across different domains.
@@
    Capturing Network Complexity. GNNs are often trained on a smaller network size and relatively smaller dataset which are insufficient to capture the underlying complexity of the network. Another possible situation with collected datasets is that they can be imbalanced. This imbalance can occur when more data points are collected for a perfect scenario without any abnormal or noisy cases. When this data is used for training a GNN model, the developed model will work well for perfect cases while failing in cases with noise or any perturbations. Techniques such as Data Augmentation [263], and Generative Adversarial Network [264] can be integrated with GNN as one of the possible solutions to address the challenge of limited data. The research community is also encouraged to explore the integration of Transfer Learning, Meta Learning as additional possible solutions.
@@
    Interpretable and Explainable GNN. One of the major challenges with any ML-based algorithms is the lack of interpretability. Interpretability can be defined as the measure of how much the user can understand the decision of the model. In other words, if the user can understand and predict the output of the ML algorithm then the ML algorithm/model is said to be highly interpretable. Similarly, the users should also be able to understand the factors that affect the performance of the GNN model and ensure that the decisions align with human intuition and domain knowledge. However, in the wireless domain (specifically IoT), with more nodes and heterogeneous data, it is challenging to design a highly interpretable and explainable GNN model. A good starting point for research along the design of interpretable and explainable GNN is GNNExplainer [265], the first-general, model-agnostic approach that can provide interpretable explanations for any GNN-based model on any graph-based learning tasks. This line of research is very important to further enhance GNN’s features, making it a reliable and preferred ML technique across many domains.
@@
    Preservation of Privacy and Security. The current applications of GNN for IoT and wireless networks collect data from different sensors and users as well as data shared between sensors (or users). While this is not an area of concern for simulation-based implementations for proof-of-concepts implementations, it is important to integrate privacy and security measures into the GNN algorithm. One of the most promising and well know techniques for preserving privacy and security is Federated Learning (FL). In FL, instead of sharing the data between nodes/sensors, only the model weights are shared. This enables collaborative learning without compromising data privacy and security. An example of GNN using FL is FedGraphNN[266] which is an open FL system for research on Federated GNNs. The integration of GNN with Federated Learning is an interesting and important area of research for future IoT and next-generation wireless networks.
@@
    Zero-Touch GNN. In the current implementation of GNNs, the algorithm designer has to grapple with the design of GNN making the process tedious for complex network scenarios. Automated Machine Learning (AutoML) [267] promises to be a solution to overcome the tedious design process. AutoML is an enabling technique to improve the implementation of ML models such as GNN by automatically learning the important steps of the ML model design process. There is some progress towards Zero-Touch GNN such as Auto-GNN [268] which helps in finding the optimal GNN architecture within a predefined search space and Auto-STGCN [269] which helps in generating optimal STGCN models automatically by quickly scanning over the parameter search space. Zero-Touch GNN will play an important role in increasing the efficiency of the GNN modeling process and requires significant attention from the research community to make it a widely adopted strategy in designing the GNN model for the future of wireless networks.
@@
    Bridging the Sim2Real Gap. In most of the existing literature, the researchers tend to focus more on two aspects of validating the performance of designed the GNN model. First is the use of open-source datasets and second is the simulation environment. This creates a gap between the performance of the proposed GNN algorithm in a simulation environment and the real-world environment. There are several National Science Foundation (NSF) Platforms for Advanced Wireless Research (PAWR) platforms such as Colosseum [270], Platform for Open Wireless Data-driven Experimental Research (POWDER) [271], Cloud Enhanced Open Software Defined Mobile Wireless Testbed for City-Scale Deployment (COSMOS) [272], Agriculture and Rural Communities (ARA) [273], Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) [274] and other testbeds such as ARENA [275], CloudRAFT [276], UnionLabs [277]. All these platforms are available to the research community to conduct real-world experiments. The researchers can share the test scenario, code, and dataset with the community to accelerate the growth of GNN as well as next-generation wireless networks.
@@
    Cross-Layer Optimization. In wireless networks, each layer is interconnected, and changes made in one layer have direct or indirect effect on other layers. Currently, most of the state-of-the-art literature of GNN focuses on optimizing only one layer, for example, the physical layer sensing or network layer routing. However, it is important to jointly optimize more than one layer to achieve better network performance [278, 279]. A rare example of GNN-based cross-layer optimization is GCLR [280] which is a GNN-based cross-layer optimization system for Multipath TCP protocol. This framework helps in predicting the expected throughput and the best routing path based on the given network topology and multipath route candidate set. While traditional optimization has been extensively used for cross-layer optimization problems [281, 282, 283, 284, 285, 286], GNN promises to be one of the best solutions for large-scale, dynamically changing networks.
@@ VIII Conclusion
In this article, we presented a comprehensive survey of GNN with a focus on the IoT and NextG networks. First, we provided a detailed description of GNN by starting with the common terminologies and their definitions before elaborating on its architecture. To complete this introduction to GNN, we provide detailed insight into the commonly used GNN variants. Second, we discussed the applications of GNN for IoT where we reviewed the existing literature focusing on IoT Network Intrusion Detection and IoT Data Fusion. Third, we looked into how GNNs have been leveraged for spectrum awareness and reviewed the existing literature focusing on RF signal classification and RF spectrum sensing. Fourth, we delved into applications of GNN for networking. Here, we discussed the existing literature focusing on applications of GNN for network and signal characteristics prediction, routing optimization, congestion control, mobile edge computing, digital twin networks, and unmanned aerial networks. Fifth, we focused on applications of GNN for tactical systems where we discussed existing literature on tactical communication and sensing systems, target recognition and localization. Finally, based on the detailed survey that we have conducted, we have listed the possible research directions for improving GNN features as well as open research areas that will benefit from GNN’s versatility.

### v2

@@ VII Opportunities and Roadmap ahead
As seen in this article, GNNs have been widely adopted for a wide range of applications but several areas are in search of breakthroughs to make GNNs more impactful. This provides an opportunity for the research community to further develop new features for GNN as well as apply GNN for even broader research areas. To this end, in this section, we discuss the possible avenues of research to improve the capabilities of GNN and research areas where GNN will play a crucial role in IoT and NextG networks. Figure 17 shows the key areas of research opportunities.
@@
    Scalable and Reliable GNN. Scalability is one of the crucial aspects to gauge the success of any ML algorithm and GNN is not an exception. As of now, most of the GNN algorithms are tested on a smaller network topology and present good results with a slightly larger network to show the scalability. However, with the evolving number of IoT devices and wireless devices, it is crucial to develop algorithms that can scale up to thousands or hundreds of thousands of devices without any possible degradation in performance. GNN promises to be a great tool that can enable scalable ML models, it is important for the research community to further investigate and rigorously test their algorithm from the perspective of scalability on real networks. Reliability is another important area that requires further attention. As of now, most of the GNN-based algorithms designed assume a perfect environment with pre-defined interference, channel, and other relevant models for system design. It is important to design GNNs that are robust to adversarial attacks, sensor malfunction, data corruption, and other unforeseen conditions. This is especially important when there are millions of connected devices and data from one or more sensors should not impact the whole network. This can be addressed by systematically identifying, quantifying, and rectifying anomalies or security risks. This promises to be an interesting area of research and will significantly benefit in maturation and adoption of GNN across different domains.
@@
    Capturing Network Complexity. GNNs are often trained on a smaller network size and relatively smaller dataset which are insufficient to capture the underlying complexity of the network. Another possible situation with collected datasets is that they can be imbalanced. This imbalance can occur when more data points are collected for a perfect scenario without any abnormal or noisy cases. When this data is used for training a GNN model, the developed model will work well for perfect cases while failing in cases with noise or any perturbations. Techniques such as Data Augmentation [340], and Generative Adversarial Network [341] can be integrated with GNN as one of the possible solutions to address the challenge of limited data. The research community is also encouraged to explore the integration of Transfer Learning, Meta Learning, Edge-perturbation, Motif-Sampling as additional possible solutions.
@@
    Interpretable and Explainable GNN. One of the major challenges with any ML-based algorithms is the lack of interpretability. Interpretability can be defined as the measure of how much the user can understand the decision of the model. In other words, if the user can understand and predict the output of the ML algorithm then the ML algorithm/model is said to be highly interpretable. Similarly, the users should also be able to understand the factors that affect the performance of the GNN model and ensure that the decisions align with human intuition and domain knowledge. However, in the wireless domain (specifically IoT), with more nodes and heterogeneous data, it is challenging to design a highly interpretable and explainable GNN model.A good starting point for research along the design of interpretable and explainable GNN is GNNExplainer [342], the first-general, model-agnostic approach that can provide interpretable explanations for any GNN-based model on any graph-based learning tasks. This line of research is very important to further enhance GNN’s features, making it a reliable and preferred ML technique across many domains. Another technique for explainable GNN is called Integrated-Gradients (IG) methods [343]. A version of IG was used for GNN-based segmentation networks [344] where IG is used to explain a DualGCN model applied in the image segmentation domain.
@@
    Preservation of Privacy and Security. The current applications of GNN for IoT and wireless networks collect data from different sensors and users as well as data shared between sensors (or users). While this is not an area of concern for simulation-based implementations for proof-of-concepts implementations, it is important to integrate privacy and security measures into the GNN algorithm. One of the most promising and well-known techniques for preserving privacy and security is Federated Learning (FL). In FL, instead of sharing the data between nodes/sensors, only the model weights are shared. This enables collaborative learning without compromising data privacy and security. An example of GNN using FL is FedGraphNN[345] which is an open FL system for research on Federated GNNs. The integration of GNN with Federated Learning is an interesting and important area of research for future IoT and next-generation wireless networks. There are also other techniques such as Homomorphic Encryption [346], Differential Privacy [347], that can be used to preserve the data privacy for GNN models.
@@
    Zero-Touch GNN. In the current implementation of GNNs, the algorithm designer has to grapple with the design of GNN making the process tedious for complex network scenarios. Automated Machine Learning (AutoML) [348] promises to be a solution to overcome the tedious design process. AutoML is an enabling technique to improve the implementation of ML models such as GNN by automatically learning the important steps of the ML model design process. There is some progress towards Zero-Touch GNN such as Auto-GNN [349] which helps in finding the optimal GNN architecture within a predefined search space and Auto-STGCN [350] which helps in generating optimal STGCN models automatically by quickly scanning over the parameter search space. There are also other techniques such as Graph Neural Network Neural Architecture Search (GraphNAS) [351], GraphNAS++ [352], Parallel Graph Architecture Search GraphPAS [353], Differentiable Architecture Search (DARTS) [354] that can be used to design best GNN models automatically. Zero-Touch GNN will play an important role in increasing the efficiency of the GNN modeling process and requires significant attention from the research community to make it a widely adopted strategy in designing the GNN model for the future of wireless networks.
@@
    Bridging the Sim2Real Gap. In most of the existing literature, the researchers tend to focus more on two aspects of validating the performance of designed the GNN model. First is the use of open-source datasets and second is the simulation environment. This creates a gap between the performance of the proposed GNN algorithm in a simulation environment and the real-world environment. There are several National Science Foundation (NSF) Platforms for Advanced Wireless Research (PAWR) platforms such as Colosseum [355], Platform for Open Wireless Data-driven Experimental Research (POWDER) [356], Cloud Enhanced Open Software Defined Mobile Wireless Testbed for City-Scale Deployment (COSMOS) [357], Agriculture and Rural Communities (ARA) [358], Aerial Experimentation and Research Platform for Advanced Wireless (AERPAW) [359] and other testbeds such as ARENA [360], CloudRAFT [361], UnionLabs [362]. All these platforms are available to the research community to conduct real-world experiments. In addition there are also other open-source platforms such as Open Network Automation Platform (ONAP) [363], which is a unified platform for orchestrating, managing, and automating network and edge computing services across telecom operators, cloud providers, and enterprises while Central Office Re-architected as a Datacenter (CORD) [364] transforms legacy central offices into cloud-native edge data centers using SDN/NFV to support low-latency, flexible service delivery. Furthermore, most articles reviewed in this survey fails to discuss how the proposed GNN frameworks could integrate with existing 3GPP or IEEE protocols. It is recommended that future research discuss this integration to demonstrate the possibility of adopting the proposed solution in real-world large-scale commercial networks. The researchers can share the test scenario, code, and dataset with the community to accelerate the growth of GNN as well as next-generation wireless networks.
@@
    Cross-Layer Optimization. In wireless networks, each layer is interconnected, and changes made in one layer have direct or indirect effects on other layers. Currently, most of the state-of-the-art literature of GNN focuses on optimizing only one layer, for example, the physical layer sensing or network layer routing. However, it is important to jointly optimize more than one layer to achieve better network performance [365, 366]. A rare example of GNN-based cross-layer optimization is GCLR [367] which is a GNN-based cross-layer optimization system for Multi-path TCP protocol. This framework helps in predicting the expected throughput and the best routing path based on the given network topology and multi-path route candidate set. While traditional optimization has been extensively used for cross-layer optimization problems [368, 369, 370, 371, 372, 373], GNN promises to be one of the best solutions for large-scale, dynamically changing networks.
@@ VIII Conclusion
In this article, we presented a comprehensive survey of GNN with a focus on the IoT and NextG networks. First, we provided a detailed description of GNN by starting with the common terminologies and their definitions before elaborating on its architecture. To complete this introduction to GNN, we provide detailed insight into the commonly used GNN variants. Second, we discussed the applications of GNN for IoT where we reviewed the existing literature focusing on IoT Network Intrusion Detection and IoT Data Fusion. Third, we looked into how GNNs have been leveraged for spectrum awareness and reviewed the existing literature focusing on RF signal classification and RF spectrum sensing. Fourth, we delved into applications of GNN for networking. Here, we discussed the existing literature focusing on applications of GNN for network and signal characteristics prediction, routing optimization, congestion control, mobile edge computing, digital twin networks, and unmanned aerial networks. Fifth, we focused on applications of GNN for tactical systems where we discussed existing literature on tactical communication and sensing systems, target recognition, and localization. Finally, based on the detailed survey that we have conducted, we have listed the possible research directions for improving GNN features as well as open research areas that will benefit from GNN’s versatility.